OpenStudy (anonymous):

How would you finish this? I have this 3.3x+5=-5.4 3.3x + 5-5=-5.4-5 3.3x=-10.4

OpenStudy (precal):

why do you have 2 equal signs on step 2?

OpenStudy (precal):

divide by 3.3 on both sides (last step)

OpenStudy (ganpat):

what was that ?? it should be 3.3x+5=-5.4 3.3x +5 -5=-5.4-5 3.3x = 0.4

OpenStudy (precal):